INAF LBC+MODS queue observing, 2012 February 19-20

Observer: K. Boutsia, M. Fumana
Telescope Operator: D. Gonzales-Huerta
Support Astronomer: D. Thompson
Instrument: LBC+MODS

Summary

Overcast skies at sunset, humidity is high and the wind is >18m/s. Small chances of opening.
While we are closed we start MODS calibrations.

At 10:20UT we have low humidity and clear skies. We open the dome. The seeing is really bad ~3.0''.
After collimating we slew to the first field by Gilli but we take data mostly to monitor the seeing...

No useful data acquired - seeing variable between 1.5" and 3.0"
No flats taken. Dome closed because to the wind at 12:50UT.

Data Summary

Project Objects Filter ExpTime Notes
Gilli field1 r-sloan / i-sloan ~1hr all data out of specs - seeing >1.5"

Details

00:30 10Bias_Bino

06:28 Since we are still closed - we start doing MODS calibrations
  • execMODS gmosARCNonino0647.cal mods1b 0002-0004 / mods1r 0002-0004 (ID520958) - there is no MASKNAME in the fits header but the mask has been correctly inserted
  • Executing gmosFlatNonino0647.cal mods1b 0005-0009 / mods1r 0005-0009 (ID520958, VFlat)- there is no MASKNAME in the fits header but the mask has been correctly inserted
  • Executing gmosFlatNonino0647new.cal mods1b 0010-0019 / mods1r 0010-0019 (ID520958, Qth1)- there is no MASKNAME in the fits header but the mask has been correctly inserted
  • Mods command line: blue/red fitsflush (not during readout!!)

  • Executing gmosARCNonino0744.cal mods1b 0020-0022 / mods1r 0020-0022 (ID581623)
  • Executing gmosFlatNonino0744.cal mods1b 0023-0027 / mods1r 0023-0027 (ID581623, VFlat)
  • Excecuting gmosFlatNonino0744new.cal mods1b 0028-0037 / mods1r 0028-0037 (ID581623, Qth1)

  • Excecuting grmosflats.55.FIORE3.cal mods1b 0038-0047 / mods1r 0038-0047 (ID573711, Qth1)
  • Excecuting grmoslamps.55.FIORE3.cal mods1b 0048-0050 / mods1r 0048-0050 (ID573711)
  • Excecuting grmosflats.55.FIORE3new.cal mods1b 0051-0055 / mods1r 0051-0055 (ID573711, VFlat)

  • Excecuting grmosflats.57.FIORE2.cal mods1b 0056-0065 / mods1r 0056-0065 (ID574126, Qth1)
  • Excecuting grmosflats.57.FIORE2new.cal mods1b 0066-0070 / mods1r 0066-0070 (ID574126, VFlat)
  • Excecuting grmoslamps.57.FIORE2.cal mods1b 0071-0073 / mods1r 0071-0073 (ID574126)

  • Excecuting grlamps1.0.cal mods1b 0074 / mods1r 0074 (LS5x60x1.0) - terminated with errors - REPEAT!
  • Excecuting SlitFlatLBT1.0.cal mods1b 0075-0079 / mods1r 0075-0079 (LS5x60x1.0, VFlat)

  • Excecuting grlamps5.0.cal mods1b 0080-0082 / mods1r 0080-0082 (LS60x5)
  • Excecuting SlitFlatLBT5.0.cal mods1b 0083-0087 / mods1r 0083-0087 (LS60x5, VFlat)

  • execMODS grlamps1.0.cal mods1b 0088-0090 / mods1r 0088-0090 (LS5x60x1.0)

10:20 Open the dome

10:29 Slew to focus1

10:32 dofpia,/first,/X2
  • dimm seeing 3"
-11675      0      0      0      0      0      0   2.01   4.60     red
-3258     96   -412    -58    -52    337     78   1.92   1.31    blue

-7     38    163   -650     82    378     82   2.22   0.39     red
-1336     24     24   -723    -52      0      0   3.06   0.62    blue

-676     69    128   -153      4    299     54   2.03   0.35     red
-973   -170   -257   -364   -413      0      0   3.00   0.47    blue

 -323    -82     99    124    -38    109     57   1.86   0.18     red
  No GOOD pupils found... check input image!

10:39 dofpia,/first,/X2 again
 2935    157     77    297    172     19     13   1.84   1.17     red
1056     57   -166    285   -194     83     77   2.33   0.46    blue

-1089    -95    -42    -90    188    104     21   1.56   0.45     red
-1516    -54    -50      4    267     76     38   1.66   0.62    blue

 277      0     97   -150   -180   -108      9   1.52   0.19     red
571    -21   -151    280   -156      0      0   2.52   0.28    blue

10:54 dofpia,/first,/X2 again
4365     71    175    765    210   -249    -44   2.42   1.77     red
2293    159    -70    266   -176     -6      3   1.53   0.92    blue

-303     72    121    209    191     12     22   2.11   0.20     red
 277     -7   -219    317    465      0      0   3.40   0.29    blue

10:59 Slew to ri1-1
  • seeing measured on images ~2.25
  • seeing 2.2blue, 1.9red
  • seeing 2.1 blue, 1.94 red - out of specs

11:30 dimm seeing 2.5"

11:37 dofpia,/first,/X2
 1135     48     63     73    114   -244    -37   1.35   0.48     red
 2203   -142     99    -18   -130   -263    -42   1.37   0.89    blue

 -134     69     30   -210   -276   -245    -29   1.29   0.24     red
-679   -213     78    139      0   -257    -42   1.21   0.34    blue

 739   -113    -30     79     91   -182    -26   1.30   0.33     red
 644   -105     16   -131     55   -212    -41   1.24   0.31    blue

 1056     50    -55    -53    -28   -114      3   1.37   0.43     red
 947   -119    -45     22   -135    -92      3   1.49   0.39    blue

  -669    -13   -124     58     -1   -138    -17   1.27   0.29     red
 -243    -26     66     21    179    -65     -5   2.08   0.16    blue

11:50 repeat ri1-1
  • previous set was out of spec and now seeing measurements have improved a bit
  • seeing measured on blue:2.0" and on red:1.6"
  • dimm seeing 1.5"
  • seeing measured on blue:2.8" and on red:2.0"
  • seeing measured on blue:2.8" and on red:1.7"
  • seeing measured on blue:2.25" and on red:1.6"

12:28 dofpia,/first,/X2
 -304     73     47    -17   -310    -83      4   1.30   0.21     red
 3177    134    -55    -67    223    -68     -4   1.83   1.26    blue

1000     45    -42   -235   -111   -257    -23   1.13   0.45     red
 155   -159    100      0    214   -152    -13   1.17   0.18    blue

12:33 Seeing too high. Observations aborted. Waiting some minute to acquire flats

12:50 Dome close because of wind.

-- KonstantinaBoutsia - 20 Feb 2012
Topic revision: r3 - 20 Feb 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback